OpenStudy (anonymous):

Evaluate 1 + 3 + 6 + 10 + 15 + ... + 500500.

OpenStudy (anonymous):

Medals will be given to anyone that has a different solution than anyone else's.

OpenStudy (anonymous):

ok this seems like an interesting one...

OpenStudy (anonymous):

this looks like the sequence (n^2 +n)/2 from n=1 to n=1000

OpenStudy (anonymous):

good,and so... ? What is the answer? That's a good method!

OpenStudy (anonymous):

xactxx the best way to do these kind of problems is to program them on the computer its not easy to do donkey work once you get series it's really easy to come up with an easy program in C or other languages..

OpenStudy (anonymous):

1+1(x+2)+...+500500

OpenStudy (anonymous):

not really, there's a perfectly simple solution without using a programming language.

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

just curious what class is this for? you seem to be doing a lot of series problems..

OpenStudy (anonymous):

no class at all, just posed this question to give medals to people.

OpenStudy (anonymous):

the sum of of the series x^2 from x=1 to n is n(n+1)(2n+1)/6 so 1000(1001)(2001)/6 = 333,833,500

OpenStudy (anonymous):

x=1 then x=2 then x=3...

OpenStudy (anonymous):

x=1 then x=2 then x=3...

OpenStudy (anonymous):

now divide that b 2 and you get 166,916,750

OpenStudy (anonymous):

now, add the sum of the series 1,2,3...n from 1 to 1000 which is (n+1)(n/2) and you get 500,500. divide that by 2 and you get 250,250. add the 2 numbers and you get 167,167,000

OpenStudy (anonymous):

i hope i didn't make any mistakes with this one...

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

wow thanks for the medals guys :) you know of any more interesting problems? i tend to get bored easily lol